calculate Ka for 0.10 M NH4Cl (pH= 4.86) ?
Answer #1

pH = 4.86

pH = -log [H+]

4.86 = -log [H+]

[H+] = 1.38 \times 10-5 M

NH4+\rightarrow NH3 + H+
The equilibrium constant for this weak acid is:
Ka = [NH3] [H+] / [NH4+]

Using ICE, we would get for the equilibrium (i.e. final) concentrations (let x = [H+]):

[NH4+] = 0.1 - x (initial 0.1 M, change -x)
[NH3] = x (initial 0, change + x)
[H+] = x (initial 0, change + x)

Ka = x2 / 0.1-x

0.1>>>x, so we can write 0.1 instead of 0.1-x

Ka = x2 / 0.1 = (1.38 \times 10-5)2/0.1 = 1.90 \times 10-9